OpenStudy (anonymous):

Using the concept of slope, find t such that the three points are collinear: (-3,3), (t,-1),(8,6)

OpenStudy (anonymous):

t=-53/3

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

To find the slope you use:\[m=(y _{2}-y _{1})/(x _{2}-x _{1})\] Using the ordered pairs you have: (6-3/8-3)=3/5

OpenStudy (anonymous):

Now we find the intercept using a point. y=mx+b; (3/5)3+b=3 9+5b=15 b=6/5

OpenStudy (anonymous):

Sorry, I missed a (-) sign in the problem:slope should be: (6-3)/(8-(-3))=3/11!

OpenStudy (amistre64):

x:8 as 7:3 8(7)/3 = 56/3 ... maybe? lol

OpenStudy (anonymous):

So (3/11)(3) +b = 3 ?

OpenStudy (anonymous):

you're losing the sign too! b then becomes y=mx+b; (3/11)(-3)+b=3 b=42/11

OpenStudy (anonymous):

(3/11)t+(4/11)2=-1 t=-53/3

OpenStudy (anonymous):

Sorry for the detour!

OpenStudy (amistre64):

(-3,3) (t,-1) (8,6) +3-3 +3-3 +3-3 ------------------ (0,0) <t+3,-4> <11,3> -3/4<t+3,-4>= <11,3> -3t-9 ------ = 11 4 -3t-9 = 44 -3t = 44+9 t = -53/3 .... better

OpenStudy (amistre64):

i forgot how to subtract 6-3 :)

OpenStudy (anonymous):

Amistre how did you get -3/4 on the 4th line?

OpenStudy (anonymous):

OH nvm i think i get it now thanx for explaining guys :D

OpenStudy (amistre64):

i needed my vectors to match; so a scaled the unknown my -3/4